Question text
A 35.4 g bullet strikes a 18.0 kg stationary wooden block and embeds itself in the block. The block and the bullet fly off together at 13.3 m/s.



What was the initial velocity of the bullet?
1
Expert's answer
2020-04-16T09:02:18-0400


We write the law of conservation of momentum in the projection onto the axis X

"m \cdot u=(m+M) \cdot V"

where will we write

"u= \frac{(m+M)}{m} \cdot V=\frac{(0.0354+18)}{0.0354} \cdot 13.3=6776 m/s=6.776 km/s"